# A Microscopic Perspective on Moiré Materials

### v1

@@ I.1 Six Years of Moiré Materials
Today, there are roughly a dozen distinct moiré material platforms designed by this prescription that have been the subject of immense theoretical and experimental research efforts. The rotational misalignment design motif describes a wide variety of moiré homobilayers, including twisted monolayer graphene (ex. magic-angle twisted bilayer graphene [4, 5, 6, 7], Fig. 1b), twisted multilayer graphene (ex. twisted monolayer-bilayer graphene [8, 9, 10], twisted double bilayer graphene [11, 12, 13]) and twisted transition metal dichalcogenides (ex. twisted bilayer WSe [14, 15, 16], WS [17, 18], WTe [19, 20] and MoTe [21, 22, 23]; Fig. 1c), and has further expanded to include multilayer structures with multiple twisted interfaces (ex. magic-angle twisted trilayer / quadralayer / pentalayer graphene [24, 25, 26, 27], moiré quasicrystalline trilayer graphene [28], helical trilayer graphene [29]). The lattice mismatch motif describes a series of aligned heterobilayer transition metal dichalcogenides (ex. aligned WSe / WS [30, 31], Fig. 1e; aligned MoTe / WSe [32, 33, 34], Fig. 1f) and aligned graphene / hexagonal boron nitride (hBN) heterostructures (ex. monolayer / bilayer graphene aligned to hBN [35, 36, 37, 38, 39]). Furthermore, many more moiré material platforms have been predicted to host exotic correlated electronic states, and are awaiting experimental realization [40, 41, 42, 43, 44, 45, 46, 47, 48].
@@ I.2 Rich, Tunable Phase Diagrams of Moiré Materials
Fig. 1c represents the phase diagram of twisted bilayer MoTe (tMoTe) as a function of out-of-plane electric field and filling. Unlike MATBG, tMoTe admits an electric-field tunable flat band structure derived from the inequivalent stacking regions of its uniquely hexagonal moiré superlattice. tMoTe displays spontaneous orbital ferromagnetism within a wide range of fillings [21], which additionally stabilizes a quantum anomalous Hall (QAH) insulator (i.e. Chern insulator at zero magnetic field) at filling and fractional Chern insulators (FCIs) showing fractional quantum anomalous Hall effects at fillings [22, 23, 70, 71]. A series of topological phase transitions are induced with increasing electric field, resulting in correlated insulating phases at these fillings. These are likely Mott and generalized Wigner crystal (GWC) states [31, 30], although verification of this hypothesis will require further investigation using local imaging techniques. Finally, a compressible state with a large Hall resistance of 2h/e appears near filling [70], possibly an anomalous composite Fermi liquid regime reminiscent of the half filling state of the lowest Landau level [73, 74].
Fig. 1e represents the phase diagram of aligned WSe / WS as a function of temperature and filling. The moiré superlattice potential of heterobilayer moiré materials made from TMDs is deeper than those of twisted graphene-based moiré materials [75], which favors charge-ordered states where electrons / holes are trapped near high-symmetry-stacking locations of the moiré superlattice. In aligned WSe / WS , charge-ordered Mott insulators are observed near [30, 31], and GWC phases (i.e. commensurate filling electron solid phases) are observed near roughly two dozen fractional fillings [31, 76, 77]. These charge-ordered phases will be the focus of Section IV.
Finally, Fig. 1f represents the phase diagram of AB-stacking-aligned MoTe / WSe as a function of out-of-plane electric / displacement field and filling. Here, “AB-stacking” refers to a near twist angle between MoTe and WSe layers, distinct from “AA-stacking” (twisted near ) because of the threefold rotational symmetry of the TMD layers. Near , AB-stacked MoTe / WSe hosts a correlated insulator, the nature of which depends on the magnitude of the electric field [32, 34]. The Mott insulator is doped into a metallic state with quantum oscillations associated with the bands of both MoTe and WSe. The charge-transfer insulator (CT Ins.) is doped into a metal with quantum oscillations only associated with the WSe layer, while the MoTe layer remains insulating and hosts a lattice of local moments. The coupling between these layers produces a tunable moiré Kondo lattice regime [34], persisting to , where a quantum spin Hall (QSH) insulator appears over a broad range of displacement field [33]. Further increasing the displacement field at produces a Chern insulator with a quantized anomalous Hall response [32].
@@ I.3 Capabilities of Local Probe Techniques
Voltage bias the sample and measure the quantum tunneling current through the tip-sample junction. Measures local density of states (LDOS ) with eV energy resolution and atomic-scale spatial resolution. Limited to back-gated device geometries. |
Spectroscopically measure flat bands and their one-particle excitations (i.e. spectral function) Density-dependent spectroscopy of energetically gapped phases (i.e. insulating, superconducting, topological phases) Wavefunction mapping of correlated phases in ultra-clean materials |
Measure the 2-terminal conductance of the tip SET, using a local gate to maintain voltage sensitivity. Measures electronic compressibility () with nm spatial resolution and 10 V/Hz voltage sensitivity. Limited to back-gated device geometries (but compatible with hBN encapsulation). |
Measure the chemical potential () and electronic compressibility () of many-body phases Identify thermodynamic phase transitions between correlated phases Measure the entropy (S), electric polarization (), and magnetization () of correlated phases via T-, -, and -dependence, respectively |
Shine monochromatic light on the sample and measure the energy and momentum of ejected electrons. Measures energy-momentum dispersions (DOS(E,k)) with meV energy resolution and sub-m spatial resolution. Limited to back-gated device geometries and elevated temperatures (T K). |
Spectroscopically visualize the dispersion characteristics of the flat bands themselves Identify collective excitations that strongly couple to electronic degrees of freedom (ex. phonons, plasmons, magnons, etc.) |
@@
As a thermodynamic probe, scanning SET gives access to the chemical potential and the electronic compressibility of moiré materials [86, 68, 87, 69]. This technique is singular in its ability to identify thermodynamic phase transitions between correlated phases, and to do so in microscopic ( m) regions of devices where fragile quantum phases can be stabilized. In addition, by performing such measurements as a function of temperature, out-of-plane electric field, or out-of-plane magnetic field (in conjunction with Maxwell’s relations), researchers gain access to the entropy, the electric polarization, or the magnetization of the system, respectively [88]. Such information can not only complement transport and spectroscopic experiments, but also provides insights into the nature of complex correlated phases that are otherwise inaccessible.
Lastly, many electromagnetic probes are used to investigate correlated phases that admit broken inversion-symmetry and broken time-reversal symmetry, which produce measurable electrostatic and magnetostatic responses, respectively. Nano-SQUID techniques have characterized the magnetic response of orbital magnetic phases ubiquitously found in moiré materials [89, 90]. Scanning NV magnetometry has identified moiré magnetism, spatially modulating spin configurations, in twisted CrI [91, 92]. KPFM, EFM, and SNOM measurements have identified ferroelectric phases and distinctive lattice dynamics occurring in stacking-engineered polar moiré materials [93, 94, 95, 96, 97]. Overall, the spatial resolution of these techniques provides unique insights into the domain structure and electric domain-switching mechanisms of magnetic and ferroelectric phases in moiré materials.
@@ I.4 Advantages of Local Probe Techniques
Finally, local probe techniques are particularly informative in the presence of intrinsic (i.e. electronic) inhomogeneity. This advantage helps distinguish the microscopic mechanisms responsible for electronic phases, and emphasizes the importance of spatial resolution to local probe techniques. Consider the situation where electrical transport experiments measure an insulating phase in a 2D material. For that same state, local probes like the STM can use the spatial distribution of the local density of states to identify the nature of this insulating phase (ex. a charge-density wave state in NbSe [100, 101], a Mott insulating state in TaSe [102], or an incommensurate Kekulé spiral state in MATBG [84, 103, 104]). Recent innovations in van der Waals device geometries have further equipped local probe techniques in their abilities to measure new physical observables, like the real-space distribution of charge, which has enabled the direct visualization of new electronic order (ex. a series of GWC and stripe phases in aligned WSe / WS ) [105, 17].
@@ II.1 Flat Electronic Bands
This early spectroscopic study was followed by the seminal work of Rafi Bistritzer and Allan MacDonald [50], predicting the emergence of flat moiré bands at an infinite sequence of “magic angles” in twisted bilayer graphene, which contradicted the contemporary expectation that the bandwidth of moiré bands would decrease monotonically with twist angle. Instead, a delicate balance of the intralayer hopping terms and the out-of-plane tunneling terms in the Hamiltonian of twisted bilayer graphene produces flat bands at the magic angle of 1.1, away from the zero twist angle limit. As the angle is further reduced from this magic angle, the bandwidth of these lowest energy bands increases.
@@ IV.2 Charge-Ordered Phases in Moiré Transition Metal Dichalcogenides
In this Section, we highlight recent local probe experiments using microwave impedance microscopy (MIM) and scanning tunneling microscopy (STM), which probe charge-ordered phases in WSe / WS moiré superlattices while circumventing the complications of ohmic contact engineering. These experiments uncover an abundance of GWCs at nearly two dozen fractional commensurate fillings of the moiré superlattice, and even provide spatial signatures of these phases that directly confirm their translation and rotational symmetry-broken nature.
@@ IV.3 Imaging Correlated Insulators in Moiré TMDs
The earliest reports of charge-ordered states in TMD-based moiré materials investigated WSe / WS moiré superlattices, primarily using optical probes [30, 31]. These reports uncovered a correlated insulator at filling , whose integer charge occupancy per moiré unit cell was suggestive of a Mott insulator (upper left panel; Fig. 4f). This was further evidenced by signatures of local, antiferromagnetically ordered moments at , as expected in the triangular lattice Hubbard model [30]. Additionally, correlated insulators were observed near fractional fillings and [31]. The simple fractional values of charge occupancy per moiré unit cell of these insulators was suggestive of GWCs (upper right and lower left panels; Fig. 4f).
@@
Fig. 4d shows the MIM-Im signal of this experiment as a function of back gate voltage. Deep suppressions of the MIM-Im signal act as signatures of correlated insulators that appear localized to simple integer and rational filling factors of the flat band of WSe / WS [77]. These deep suppressions appeared at nearly the same gate voltages at several locations in their device. In addition, temperature-dependent studies of this signature demonstrated that the insulating suppressions vanished at filling-dependent temperatures that resembled those expected from small-scale Monte Carlo simulations of the proposed GWC states.
Subsequently, STM imaging experiments provided unequivocal evidence for these Mott and GWC states by introducing a new charge-sensing method (Fig. 4e) that enabled researchers to directly visualize these correlated ground states in real-space [105]. In these experiments, aligned WSe / WS was encapsulated from the top by a thin hBN spacer layer, covered by a monolayer graphene sensing layer that was exposed to the STM tip. When WSe / WS is electrostatically gated into a correlated insulating state, discrete charges become localized to high-symmetry-stacking regions of the moiré superlattice (Fig. 4f). Each of these trapped charges can be locally discharged by a voltage-biased STM tip, and this discharging event can be measured in dI/dV spectra of the top graphene sensing layer. Thus, this technique provides a direct measurement of the moiré superlattice area per trapped electron / hole (not necessarily the state’s Wigner-Seitz unit cell), as schematically depicted in the panels of Fig. 4f by dark outlined polygons.
@@ V.4 Correlation-Driven Topological Phases in MATBG
Correlation effects also produce fundamentally new kinds of topological phases without single-particle analogues, including fractional Chern insulators (FCIs) measured in recent scanning SET experiments on MATBG / hBN [69]. Electronic compressibility measurements confirm a quantum anomalous Hall state near [59, 61, 87], further confirming that the density range between and is well-described by a single, isolated Chern subband. At low magnetic fields (B T), calculations show that the Berry curvature of this Chern subband is concentrated in a hot-spot near the point of the moiré Brillouin zone. This Berry curvature distribution favors charge density wave states (Fig. 5a and green lines in Fig. 5g) that break moiré-translation symmetry. At moderate magnetic fields (B T), calculations show a redistribution of Berry curvature that favors FCIs (Fig. 5d and red lines in Fig. 5g), which break T-symmetry and host non-trivial topological order. Between and , FCIs are observed in MATBG / hBN with (from ) and (from ). Very recently, FCI states have been observed at zero magnetic field in two new moiré materials. These include and FCIs in tMoTe (Fig. 1c) [22, 23, 70, 71], and , , , , , and FCIs in rhombohedral pentalayer graphene aligned to hBN [153], all of which satisfy the relation . Such zero-field FCIs closely parallel the field-stabilized FCIs in MATBG / hBN, all occurring at partial fillings of a single isolated Chern subband.
@@ VI.2 Imaging Orbital Ferromagnetism
Time-reversal symmetry breaking leads to ferromagnetism, of which there are broadly two types: spin and orbital ferromagnetism. The fundamental unit of magnetism in a spin ferromagnet is the spin of an electron, with an intrinsic magnetic dipole moment of . Spin ferromagnets are common among materials containing atoms with unpaired spins, like iron, nickel, cobalt, and many rare-earth elements. Nearly all known ferromagnets are spin ferromagnets, deriving their macroscopic magnetization largely from the alignment of these spins (ex. 2D ferromagnets like CrX (X = Cl, Br, I)). Spin-orbit coupling can cause a small amount of the material’s spin ferromagnetism to mix into the orbital degree of freedom.

### v2

@@ I.1 Six Years of Moiré Materials
Today, there are roughly a dozen distinct moiré material platforms designed by this prescription that have been the subject of immense theoretical and experimental research efforts. The rotational misalignment design motif describes a wide variety of moiré homobilayers, including twisted monolayer graphene (ex. magic-angle twisted bilayer graphene [4, 5, 6, 7], Fig. 1b), twisted multilayer graphene (ex. twisted monolayer-bilayer graphene [8, 9, 10], twisted double bilayer graphene [11, 12, 13]) and twisted transition metal dichalcogenides (ex. twisted bilayer WSe2 [14, 15, 16], WS2 [17, 18], WTe2 [19, 20] and MoTe2 [21, 22, 23]; Fig. 1c), and has further expanded to include multilayer structures with multiple twisted interfaces (ex. magic-angle twisted trilayer / quadralayer / pentalayer graphene [24, 25, 26, 27], moiré quasicrystalline trilayer graphene [28], helical trilayer graphene [29]). The lattice mismatch motif describes a series of aligned heterobilayer transition metal dichalcogenides (ex. aligned WSe2 / WS2 [30, 31], Fig. 1e; aligned MoTe2 / WSe2 [32, 33, 34], Fig. 1f) and aligned graphene / hexagonal boron nitride (hBN) heterostructures (ex. monolayer / bilayer graphene aligned to hBN [35, 36, 37, 38, 39]). Furthermore, many more moiré material platforms have been predicted to host exotic correlated electronic states, and are awaiting experimental realization [40, 41, 42, 43, 44, 45, 46, 47, 48].
@@ I.2 Rich, Tunable Phase Diagrams of Moiré Materials
Fig. 1c represents the phase diagram of twisted bilayer MoTe2 (tMoTe2) as a function of out-of-plane electric field and filling. Unlike MATBG, tMoTe2 admits an electric-field tunable flat band structure derived from the inequivalent stacking regions of its uniquely hexagonal moiré superlattice. tMoTe2 displays spontaneous orbital ferromagnetism within a wide range of fillings [21], which additionally stabilizes a quantum anomalous Hall (QAH) insulator (i.e. Chern insulator at zero magnetic field) at filling and fractional Chern insulators (FCIs) showing fractional quantum anomalous Hall effects at fillings [22, 23, 70, 71]. A series of topological phase transitions are induced with increasing electric field, resulting in correlated insulating phases at these fillings. These are likely Mott and generalized Wigner crystal (GWC) states [31, 30], although verification of this hypothesis will require further investigation using local imaging techniques. Finally, a compressible state with a large Hall resistance of 2h/e2 appears near filling [70], possibly an anomalous composite Fermi liquid regime reminiscent of the half filling state of the lowest Landau level [73, 74].
Fig. 1e represents the phase diagram of aligned WSe2 / WS2 as a function of temperature and filling. The moiré superlattice potential of heterobilayer moiré materials made from TMDs is deeper than those of twisted graphene-based moiré materials [75], which favors charge-ordered states where electrons / holes are trapped near high-symmetry-stacking locations of the moiré superlattice. In aligned WSe2 / WS2 , charge-ordered Mott insulators are observed near [30, 31], and GWC phases (i.e. commensurate filling electron solid phases) are observed near roughly two dozen fractional fillings [31, 76, 77]. These charge-ordered phases will be the focus of Section IV.
Finally, Fig. 1f represents the phase diagram of AB-stacking-aligned MoTe2 / WSe2 as a function of out-of-plane electric / displacement field and filling. Here, “AB-stacking” refers to a near twist angle between MoTe2 and WSe2 layers, distinct from “AA-stacking” (twisted near ) because of the threefold rotational symmetry of the TMD layers. Near , AB-stacked MoTe2 / WSe2 hosts a correlated insulator, the nature of which depends on the magnitude of the electric field [32, 34]. The Mott insulator is doped into a metallic state with quantum oscillations associated with the bands of both MoTe2 and WSe2. The charge-transfer insulator (CT Ins.) is doped into a metal with quantum oscillations only associated with the WSe2 layer, while the MoTe2 layer remains insulating and hosts a lattice of local moments. The coupling between these layers produces a tunable moiré Kondo lattice regime [34], persisting to , where a quantum spin Hall (QSH) insulator appears over a broad range of displacement field [33]. Further increasing the displacement field at produces a Chern insulator with a quantized anomalous Hall response [32].
@@ I.3 Capabilities of Local Probe Techniques
Voltage bias the sample and measure the quantum tunneling current through the tip-sample junction. Measures local density of states (LDOS ) with eV energy resolution and atomic-scale spatial resolution. Limited to back-gated device geometries. | Spectroscopically measure flat bands and their one-particle excitations (i.e. spectral function) Density-dependent spectroscopy of energetically gapped phases (i.e. insulating, superconducting, topological phases) Wavefunction mapping of correlated phases in ultra-clean materials |
Measure the 2-terminal conductance of the tip SET, using a local gate to maintain voltage sensitivity. Measures electronic compressibility () with nm spatial resolution and 10-5 V/Hz voltage sensitivity. Limited to back-gated device geometries (but compatible with hBN encapsulation). | Measure the chemical potential () and electronic compressibility () of many-body phases Identify thermodynamic phase transitions between correlated phases Measure the entropy (S), electric polarization (), and magnetization () of correlated phases via T-, -, and -dependence, respectively |
Shine monochromatic light on the sample and measure the energy and momentum of ejected electrons. Measures energy-momentum dispersions (DOS(E,k)) with meV energy resolution and sub-m spatial resolution. Limited to back-gated device geometries and elevated temperatures (T K). | Spectroscopically visualize the dispersion characteristics of the flat bands themselves Identify collective excitations that strongly couple to electronic degrees of freedom (ex. phonons, plasmons, magnons, etc.) |
@@
As a thermodynamic probe, scanning SET gives access to the chemical potential and the electronic compressibility of moiré materials [86, 68, 87, 69]. This technique is singular in its ability to identify thermodynamic phase transitions between correlated phases, and to do so in microscopic ( m2) regions of devices where fragile quantum phases can be stabilized. In addition, by performing such measurements as a function of temperature, out-of-plane electric field, or out-of-plane magnetic field (in conjunction with Maxwell’s relations), researchers gain access to the entropy, the electric polarization, or the magnetization of the system, respectively [88]. Such information can not only complement transport and spectroscopic experiments, but also provides insights into the nature of complex correlated phases that are otherwise inaccessible.
Lastly, many electromagnetic probes are used to investigate correlated phases that admit broken inversion-symmetry and broken time-reversal symmetry, which produce measurable electrostatic and magnetostatic responses, respectively. Nano-SQUID techniques have characterized the magnetic response of orbital magnetic phases ubiquitously found in moiré materials [89, 90]. Scanning NV magnetometry has identified moiré magnetism, spatially modulating spin configurations, in twisted CrI3 [91, 92]. KPFM, EFM, and SNOM measurements have identified ferroelectric phases and distinctive lattice dynamics occurring in stacking-engineered polar moiré materials [93, 94, 95, 96, 97]. Overall, the spatial resolution of these techniques provides unique insights into the domain structure and electric domain-switching mechanisms of magnetic and ferroelectric phases in moiré materials.
@@ I.4 Advantages of Local Probe Techniques
Finally, local probe techniques are particularly informative in the presence of intrinsic (i.e. electronic) inhomogeneity. This advantage helps distinguish the microscopic mechanisms responsible for electronic phases, and emphasizes the importance of spatial resolution to local probe techniques. Consider the situation where electrical transport experiments measure an insulating phase in a 2D material. For that same state, local probes like the STM can use the spatial distribution of the local density of states to identify the nature of this insulating phase (ex. a charge-density wave state in NbSe2 [100, 101], a Mott insulating state in TaSe2 [102], or an incommensurate Kekulé spiral state in MATBG [84, 103, 104]). Recent innovations in van der Waals device geometries have further equipped local probe techniques in their abilities to measure new physical observables, like the real-space distribution of charge, which has enabled the direct visualization of new electronic order (ex. a series of GWC and stripe phases in aligned WSe2 / WS2 ) [105, 17].
@@ II.1 Flat Electronic Bands
This early spectroscopic study was followed by the seminal work of Rafi Bistritzer and Allan MacDonald [50], predicting the emergence of flat moiré bands at an infinite sequence of “magic angles” in twisted bilayer graphene, which contradicted the contemporary expectation that the bandwidth of moiré bands would decrease monotonically with twist angle. Instead, a delicate balance of the intralayer hopping terms and the out-of-plane tunneling terms in the Hamiltonian of twisted bilayer graphene produces flat bands at the magic angle of 1.1∘, away from the zero twist angle limit. As the angle is further reduced from this magic angle, the bandwidth of these lowest energy bands increases.
@@ IV.2 Charge-Ordered Phases in Moiré Transition Metal Dichalcogenides
In this Section, we highlight recent local probe experiments using microwave impedance microscopy (MIM) and scanning tunneling microscopy (STM), which probe charge-ordered phases in WSe2 / WS2 moiré superlattices while circumventing the complications of ohmic contact engineering. These experiments uncover an abundance of GWCs at nearly two dozen fractional commensurate fillings of the moiré superlattice, and even provide spatial signatures of these phases that directly confirm their translation and rotational symmetry-broken nature.
@@ IV.3 Imaging Correlated Insulators in Moiré TMDs
The earliest reports of charge-ordered states in TMD-based moiré materials investigated WSe2 / WS2 moiré superlattices, primarily using optical probes [30, 31]. These reports uncovered a correlated insulator at filling , whose integer charge occupancy per moiré unit cell was suggestive of a Mott insulator (upper left panel; Fig. 4f). This was further evidenced by signatures of local, antiferromagnetically ordered moments at , as expected in the triangular lattice Hubbard model [30]. Additionally, correlated insulators were observed near fractional fillings and [31]. The simple fractional values of charge occupancy per moiré unit cell of these insulators was suggestive of GWCs (upper right and lower left panels; Fig. 4f).
@@
Fig. 4d shows the MIM-Im signal of this experiment as a function of back gate voltage. Deep suppressions of the MIM-Im signal act as signatures of correlated insulators that appear localized to simple integer and rational filling factors of the flat band of WSe2 / WS2 [77]. These deep suppressions appeared at nearly the same gate voltages at several locations in their device. In addition, temperature-dependent studies of this signature demonstrated that the insulating suppressions vanished at filling-dependent temperatures that resembled those expected from small-scale Monte Carlo simulations of the proposed GWC states.
Subsequently, STM imaging experiments provided unequivocal evidence for these Mott and GWC states by introducing a new charge-sensing method (Fig. 4e) that enabled researchers to directly visualize these correlated ground states in real-space [105]. In these experiments, aligned WSe2 / WS2 was encapsulated from the top by a thin hBN spacer layer, covered by a monolayer graphene sensing layer that was exposed to the STM tip. When WSe2 / WS2 is electrostatically gated into a correlated insulating state, discrete charges become localized to high-symmetry-stacking regions of the moiré superlattice (Fig. 4f). Each of these trapped charges can be locally discharged by a voltage-biased STM tip, and this discharging event can be measured in dI/dV spectra of the top graphene sensing layer. Thus, this technique provides a direct measurement of the moiré superlattice area per trapped electron / hole (not necessarily the state’s Wigner-Seitz unit cell), as schematically depicted in the panels of Fig. 4f by dark outlined polygons.
@@ V.4 Correlation-Driven Topological Phases in MATBG
Correlation effects also produce fundamentally new kinds of topological phases without single-particle analogues, including fractional Chern insulators (FCIs) measured in recent scanning SET experiments on MATBG / hBN [69]. Electronic compressibility measurements confirm a quantum anomalous Hall state near [59, 61, 87], further confirming that the density range between and is well-described by a single, isolated Chern subband. At low magnetic fields (B T), calculations show that the Berry curvature of this Chern subband is concentrated in a hot-spot near the point of the moiré Brillouin zone. This Berry curvature distribution favors charge density wave states (Fig. 5a and green lines in Fig. 5g) that break moiré-translation symmetry. At moderate magnetic fields (B T), calculations show a redistribution of Berry curvature that favors FCIs (Fig. 5d and red lines in Fig. 5g), which break T-symmetry and host non-trivial topological order. Between and , FCIs are observed in MATBG / hBN with (from ) and (from ). Very recently, FCI states have been observed at zero magnetic field in two new moiré materials. These include and FCIs in tMoTe2 (Fig. 1c) [22, 23, 70, 71], and , , , , , and FCIs in rhombohedral pentalayer graphene aligned to hBN [153], all of which satisfy the relation . Such zero-field FCIs closely parallel the field-stabilized FCIs in MATBG / hBN, all occurring at partial fillings of a single isolated Chern subband.
@@ VI.2 Imaging Orbital Ferromagnetism
Time-reversal symmetry breaking leads to ferromagnetism, of which there are broadly two types: spin and orbital ferromagnetism. The fundamental unit of magnetism in a spin ferromagnet is the spin of an electron, with an intrinsic magnetic dipole moment of . Spin ferromagnets are common among materials containing atoms with unpaired spins, like iron, nickel, cobalt, and many rare-earth elements. Nearly all known ferromagnets are spin ferromagnets, deriving their macroscopic magnetization largely from the alignment of these spins (ex. 2D ferromagnets like CrX3 (X = Cl, Br, I)). Spin-orbit coupling can cause a small amount of the material’s spin ferromagnetism to mix into the orbital degree of freedom.